OpenStudy (danise):

Find the real or imaginary solutions of each equation by factoring ? x^3 - 1000 = 0

OpenStudy (michele_laino):

hint:

OpenStudy (danise):

To be honest I don't know how to do the rest !!!!!

OpenStudy (michele_laino):

we have to apply the canceling law of a product

OpenStudy (michele_laino):

namely: we have to solve these 2 equations separately: x-10=0 x^2+100+10x=0

OpenStudy (danise):

Thank you so much I can finally do my homework!!!!!

OpenStudy (michele_laino):

Thank you!

OpenStudy (anonymous):

michile 's hint is not right x^3 =1000 is all your given as you write it so what same number can you multiply 3 times to =1000 10*10*10=1000 so x =10

OpenStudy (danise):

No wonder because I was thinking how did I get 100 thanks for your help